OpenStudy (anonymous):

q A line passes through the point (-3, -3) and has a slope of .1/2 What is the equation of the line?

OpenStudy (anonymous):

what is q??

OpenStudy (anonymous):

A line passes through the point (-3, -3) and has a slope of .1/2 What is the equation of the line

OpenStudy (anonymous):

do u mean slope 1/2 not .1/2

OpenStudy (anonymous):

ok then u find the equation in the form y=mx+b

OpenStudy (anonymous):

m=1/2

OpenStudy (mew55):

use the formula y-yo=m(x-x0) y+3=1/2(x+3) y+3=1/2x+3/2-3 y=1/2x-3/2

OpenStudy (anonymous):

i need the anwer correct please to complete math and graduate in December

OpenStudy (anonymous):

so plugin the points (-3,-3) in the equation y=1/2x+b to solve for b

OpenStudy (anonymous):

then u will get the equation

OpenStudy (anonymous):

okey thanks news finished have 99 for my test math thanks

OpenStudy (anonymous):

so im gruaduate

OpenStudy (mew55):

hey Bhagyashree, is the answer i put in right?

OpenStudy (anonymous):

so -3=1/2(-3)+b b=-1.5 so your equation is y=1/2x-3/2 yes your answer is correct @mew55 I have no idea what you are talking about ... "graduate"...??????? @javier24442

OpenStudy (anonymous):

okey men thanks im finish math